### v1

# Spatiotemporal Semantic V2X Framework for Cooperative Collision Prediction

Abstract
Intelligent Transportation Systems (ITS) demand real-time collision prediction to ensure road safety and reduce accident severity. Conventional approaches rely on transmitting raw video or high-dimensional sensory data from roadside units (RSUs) to vehicles, which is impractical under vehicular communication bandwidth and latency constraints. In this work, we propose a semantic V2X framework in which RSU-mounted cameras generate spatiotemporal semantic embeddings of future frames using the Video Joint Embedding Predictive Architecture (V-JEPA). To evaluate the system, we construct a digital twin of an urban traffic environment enabling the generation of diverse traffic scenarios with both safe and collision events. These embeddings of the future frame, extracted from V-JEPA, capture task-relevant traffic dynamics and are transmitted via V2X links to vehicles, where a lightweight attentive probe and classifier decode them to predict imminent collisions. By transmitting only semantic embeddings instead of raw frames, the proposed system significantly reduces communication overhead while maintaining predictive accuracy. Experimental results demonstrate that the framework with an appropriate processing method achieves a 10% F1-score improvement for collision prediction while reducing transmission requirements by four orders of magnitude compared to raw video. This validates the potential of semantic V2X communication to enable cooperative, real-time collision prediction in ITS.
I Introduction
Traffic collisions remain one of the most severe challenges in Intelligent Transportation Systems (ITS), often leading to congestion, property damage, injuries, and fatalities. Proactive collision prediction is essential to mitigate these risks by enabling early warnings and preventive actions [9]. Traditional approaches that rely on transmitting raw video [16] or high-dimensional sensory data from roadside units (RSUs) to vehicles are not feasible under V2X constraints, where both bandwidth and latency are limited. This calls for communication-efficient, task-oriented methods that transmit only the most relevant information to support predictive safety applications.
Semantic communication (SemComm) addresses this by transmitting task-relevant features instead of raw sensory data [13]. In vehicular contexts, this enables roadside units (RSUs) to extract compressed semantic representations that capture the dynamics most relevant to safety, and vehicles to interpret them for decision-making. Crucially, if these semantic representations are predictive rather than descriptive, vehicles can receive early warnings of imminent collisions rather than delayed notifications of ongoing accidents. To evaluate such systems, digital twins play a key role in generating realistic, controllable, and reproducible traffic scenarios. In this work, we employ the Quanser Interactive Labs (QLabs) platform as a digital twin of an urban traffic system. This environment provides diverse scenarios including intersections and roundabouts, where both safe and collision cases can be simulated. The digital twin thus enables systematic video-data generation for training and evaluating predictive SemComm frameworks.
Building on this foundation, we propose a semantic V2X collision prediction framework. RSUs equipped with cameras employ the Video Joint Embedding Predictive Architecture (V-JEPA) as the semantic encoder [3]. V-JEPA does not merely summarize observed frames; it predicts the embeddings of future frames, capturing the evolving motion patterns and vehicle interactions that lead to collisions. These predicted embeddings are transmitted to vehicles via V2X links, where an attentive probe and classifier act as semantic decoders, refining the embeddings to predict whether a collision is imminent. By focusing on predictive embeddings, the system moves beyond detection toward true anticipation of dangerous events. The main contributions of this work are summarized as follows:
  • We construct a high-quality video dataset using a digital twin of an urban traffic environment (QLabs). The dataset includes diverse scenarios and is labeled into two categories, collision events and safe driving, enabling systematic evaluation of predictive collision avoidance.
  • We investigate multiple post-processing techniques (heatmaps, binary masking, and hybrid methods) to enhance task-relevant features in traffic videos, and analyze their impact on the representation quality and predictive performance of the V-JEPA framework.
  • We employ V-JEPA as the semantic encoder at RSUs to extract low-dimensional spatiotemporal embeddings of future frame from video clips. These embeddings serve as semantic messages transmitted via V2X links, and are decoded by an attentive probe with cross-attention at the vehicle side to focus on high-risk regions and motion patterns for accurate collision prediction.
This work demonstrates that combining digital twins, semantic communication, and spatiotemporal embeddings can enable predictive collision avoidance in vehicular networks, achieving both high accuracy and significant communication efficiency.
II Related Works
Collision prediction and avoidance have been studied from multiple perspectives. The authors in [8] propose a deep learning-based approach using convolutional neural networks (CNNs) to analyze surveillance video for traffic congestion detection and accident monitoring, aiming to overcome the limitations of manual observation and traditional machine learning models. Similarly, [2] employs ensemble learning methods such as Random Forest, XGBoost, L-GBM, and CatBoost combined with Shapley Additive Explanations (SHAP) to predict accident severity and identify contributing factors, providing interpretable insights into road safety analysis. These works, however, largely focus on post-hoc analysis of accidents rather than proactive prediction of imminent collisions.
Big data frameworks have also been leveraged in [14], where large-scale UK traffic datasets were processed using Apache Spark with Decision Trees, Random Forest, and Naïve Bayes to predict accident severity under diverse environmental conditions. Meanwhile, [1] introduced the I3D-CONVLSTM2D model that integrates spatial and temporal features from RGB and optical flow data for real-time accident detection on edge devices, emphasizing computational efficiency. Although spatiotemporal features improve recognition, these methods detect accidents as they occur, without forecasting future risk.
Recently, the use of LLMs has been explored for traffic video analysis. In [7], the VAD-LLaMA framework combined video anomaly detection with natural language explanations, using a Long-Term Context (LTC) module and instruction-tuned Video-LLaMA to localize and explain anomalies. Similarly, [4] proposed a hybrid framework combining Bayesian Generalized Extreme Value (GEV) models with ARIMA/ARIMAX forecasting and AI-based video analytics to estimate short-term crash risks at intersections. While these approaches improve interpretability, they rely heavily on temporal correlations in data streams rather than predictive spatiotemporal embeddings tailored for proactive vehicular safety.
YOLO-based detection frameworks are also gaining traction. In [10], YOLOv8 and YOLO11x models were trained on 15,000 traffic images for real-time accident and speeding-vehicle detection, introducing persistent multi-frame verification and human-in-the-loop validation. For proactive collision avoidance, [5] integrates CARLA simulations with YOLOv7-based object detection to anticipate collisions in connected and autonomous vehicle scenarios. Similarly, [17] enhances small-object detection and near-miss prediction by combining YOLOv7 with instance segmentation, Inverse Perspective Mapping (IPM), StrongSORT tracking, and Kalman filtering. These YOLO-based methods are effective in object-level detection and trajectory forecasting but still require transmitting large amounts of visual data for centralized processing.
In contrast, this work proposes a V2X SemComm framework where RSUs generate predictive spatiotemporal embeddings using V-JEPA. Unlike CNN- or YOLO-based methods that detect current events, V-JEPA forecasts embeddings of future frames, anticipating motion patterns that may lead to collisions. Instead of transmitting raw video, embeddings are sent as semantic messages to vehicles, which decode them with a lightweight attentive probe and classifier. This design enables real-time, bandwidth-efficient collision prediction, and, to our knowledge, is the first to integrate digital twin–based data generation, predictive embeddings, and semantic V2X communication.
III Methodology
This work develops a semantic V2X framework in which RSUs equipped with surveillance cameras and V-JEPA act as semantic encoders, and vehicles act as semantic decoders for predictive collision avoidance. Unlike image-based approaches, which provide only static snapshots, video data captures both spatial and temporal dynamics, enabling richer modeling of vehicle interactions, motion patterns, and speed estimation. V-JEPA is employed at the RSU to shift predictive modeling from pixel-level forecasting to the representation space. Instead of reconstructing raw frames, V-JEPA generates compact spatiotemporal embeddings of future frames, thereby anticipating anomalous motion patterns that precede collisions. These predicted embeddings are transmitted as semantic messages over V2X links to vehicles, where an attentive probe and classifier decode them to determine whether a collision is imminent or the driving condition is safe. By transmitting embeddings rather than raw video, the system reduces both communication overhead and computational workload while preserving predictive accuracy.
Figure 1: Data Collection Using Digital-Twin (Quanser Interactive Lab)
III-A Dataset Generation
Dataset generation for the V-JEPA modules consists of two steps: (i) video collection using a digital twin environment, and (ii) post-processing to emphasize task-relevant regions.
III-A1 Video Generation using QLabs
To obtain a realistic and diverse dataset, the QLabs platform is employed as a digital twin of an urban traffic environment [15]. QLabs models intersections, roundabouts, two-way roads, pedestrians, traffic lights, and autonomous vehicles, enabling flexible and reproducible scenario creation. The Qlabs environment is programmed to generate both safe driving and collision events, observed from four RSU-mounted surveillance cameras covering different regions of the environment. Each frame is sampled every 50 ms, producing 20 FPS (Frames per second) video clips of varying lengths depending on scenario duration. This allows for reproducible traffic scene generation and controlled experimentation for predictive safety analysis. Figure1 depicts the process of video-dataset generation.
III-A2 Video Post-Processing
The collected video clips are subjected to a post-processing stage designed to emphasize task-relevant regions and simplify downstream learning. YOLOv11 [6] is employed to detect vehicles, after which three techniques are applied: (i) heatmaps, (ii) binary road masks, and (iii) a hybrid of heatmaps and binary masking. In the heatmap approach, vehicle coordinates are highlighted first, followed by road regions, while non-road areas such as surrounding trees and infrastructure are down-weighted. The binary masking approach assigns zero weight to all pixels outside the road, whereas the hybrid method further highlights vehicle coordinates while suppressing non-road regions entirely. These techniques highlight dynamic objects and road areas while reducing background redundancy with examples shown in Figure2.
Figure 2: Different Post-processing Methods for Video Dataset
III-B Semantic Encoder: V-JEPA at RSU
The V-JEPA model [3] is deployed at RSUs to generate predictive semantic embeddings. Unlike conventional predictive models, V-JEPA forecasts embeddings of masked future frame regions rather than reconstructing pixels, thereby focusing on task-relevant semantics for downstream prediction.
III-B1 V-JEPA Pretraining with Masked Modeling
As shown in Figure3, V-JEPA is pretrained in a self-supervised fashion using masked prediction. A video clip is divided into non-overlapping spatiotemporal patches. The target encoder processes the full unmasked clip, while the context encoder encodes a masked version. A predictor network generates embeddings for the masked tokens, which are trained against target encoder embeddings using an $L_{1}$ loss. This allows the model to learn predictive spatiotemporal representations in the embedding space.
Figure 3: V-JEPA Pretraining & Fine-Tuning for Collision Prediction and V2X Semantic Communication
III-B2 Fine-Tuning with Attentive Probe
For downstream collision prediction, the pretrained V-JEPA encoder takes as input a video clip of dimension $N\times W\times H$, where $N$ is the number of frames, and $W$ and $H$ denote frame width and height. The clip is flattened into a sequence of non-overlapping spatiotemporal patches, producing token embeddings of shape $N\times P\times D$, where $P$ is the number of patches per frame and $D$ is the embedding dimension. These frozen encoder parameters preserve the rich self-supervised features learned during pretraining. On top of the encoder outputs, an attentive probe applies a query-driven cross-attention mechanism to pool the tokens into a single compact vector $1\times D$, selectively aggregating information from the most relevant spatial and temporal regions. This pooled embedding is then passed to a lightweight classifier to infer one of two states: collision or safe driving. By transforming raw video data ($N\times W\times H$) into a compact embedding $1\times D$, the framework enables efficient semantic transmission while maintaining predictive accuracy for V2X communication.
III-C Semantic Decoder at Vehicle
At the vehicle side, the received embeddings are directly processed by the semantic decoder, which consists of an attentive probe and a lightweight classifier. Vehicles receive embeddings from RSUs, refine them through query-driven cross-attention, concentrating on the most critical spatial and temporal cues that may indicate anomalous motion or collision risk. The resulting compact representation is then fed into a linear classifier to produce a binary decision: collision or safe driving.
Since the decoder operates only on low-dimensional embeddings rather than raw video streams, the computational complexity of the decoder is modest (explained in Sec. III-E) This makes the design suitable for real-time inference on resource-constrained vehicular hardware, while also significantly reducing communication overhead in the V2X link. The division of computation heavy embedding extraction at the RSU and lightweight decoding at the vehicle ensures scalability and practicality for latency-sensitive V2X networks.
III-D Transmission Efficiency in V2X Semantic Communication
A major advantage of the proposed framework lies in its ability to drastically reduce transmission payloads compared to conventional raw video communication. Let each video clip contain $N$ frames of original spatial size $H_{o}\times W_{o}$ with 3 RGB channels, where $(H_{o},W_{o})$ denote the preprocessed frame dimensions. The total raw video payload can be expressed as:
$\small S_{\text{raw}}=N\cdot H_{o}\cdot W_{o}\cdot 3\quad\text{bytes}.$ (1)
In contrast, the semantic embeddings generated by the V-JEPA encoder at the RSU are significantly more compact. Each clip produces embeddings of size $N\times P\times D$, where $P$ is the number of spatiotemporal patches per frame and $D$ is the embedding dimension. The corresponding payload is:
$\small S_{\text{sem}}=N\cdot P\cdot D\cdot b\quad\text{bytes},$ (2)
where $b$ denotes the number of bytes per element ($b=4$ for FP32, $b=2$ for FP16, $b=1$ for INT8). The semantic compression ratio is then: $\mathcal{R}={S_{\text{raw}}}/{S_{\text{sem}}}$. Assuming $N=64$, $H_{o}=W_{o}=2048$, and a post-processed input of $384\times 384$ with patch size $16\times 16$ ($P=\tfrac{384}{16}\times\tfrac{384}{16}=576$) and embedding dimension $D=1280$, the corresponding payloads are:
  • Raw video (RGB): 0.81 GB per clip.
  • Embeddings (FP32): 0.74 MB ($\mathcal{R}\approx 1095\times$).
  • Embeddings (FP16): 0.37 MB ($\mathcal{R}\approx 2190\times$).
  • Embeddings (INT8): 0.18 MB ($\mathcal{R}\approx 4380\times$).
These results demonstrate that transmitting semantic embeddings instead of raw video frames achieves up to four orders of magnitude reduction in communication payload. This enables low-latency, bandwidth-efficient semantic message exchange between RSUs and vehicles, ensuring timely and reliable collision prediction in V2X networks.
III-E Complexity Analysis of Frozen Encoder with Attentive Probe
Consider a video clip composed of $N$ frames, each of spatial size $H\times W$. The spatiotemporal tokenizer divides the clip into non-overlapping patches of size $(t_{p}\times p\times p)$, where $t_{p}$ is the temporal stride and $p$ is the spatial stride. This process yields a token sequence of length, $L=\frac{N}{t_{p}}\cdot\frac{H}{p}\cdot\frac{W}{p}$, where $L$ denotes the total number of spatiotemporal tokens fed into the encoder.
III-E1 Frozen Encoder Cost
Let the encoder embedding dimension be $D$, the number of Transformer blocks be $L_{e}$, and the multi-layer perceptron (MLP) feed-forward expansion ratio be $r$. Each Transformer block consists of a multi-head self-attention (MSA) layer followed by an MLP. The floating-point operations (FLOPs) per block can be approximated as $F_{\text{block}}=2L^{2}D+(4+2r)LD^{2}$, where the first term corresponds to attention computations ($\mathcal{O}(L^{2}D)$) and the second term represents the projection and feed-forward operations ($\mathcal{O}(LD^{2})$). The total cost of frozen encoder over all $L_{e}$ layers is therefore:
$\small F_{\text{enc}}=L_{e}\bigl(2L^{2}D+(4+2r)LD^{2}\bigr).$ (3)
During inference, the peak activation memory of the encoder (without gradient storage) scales as: $\mathcal{M}_{\text{enc}}=\mathcal{O}(LD)$, dominated by token embeddings and attention buffers.
III-E2 Attentive Probe and Classifier Cost
The attentive probe aggregates the encoder’s token sequence using a single-query cross-attention mechanism, followed by a small MLP and a linear classification layer. The computational cost of the probe can be expressed as:
$\small F_{\text{probe}}=3LD^{2}+2LD+3D^{2}+DC,$ (4)
where $C$ denotes the number of output classes (e.g., $C=2$ for collision and safe driving). However, in practice, the encoder output is frozen and the probe operates with a single query vector rather than performing full self-attention over all $L$ tokens. Consequently, the cost associated with key and value projections ($\mathcal{O}(LD^{2})$) is computed once during encoding and not repeated during probing. The probe therefore performs only lightweight attention ($\mathcal{O}(LD)$) and projection operations ($\mathcal{O}(D^{2})$), resulting in an overall complexity that scales as:
$\small\mathcal{O}(D^{2}+DC),\small$ (5)
where the $D^{2}$ term arises from the linear transformations within the probe and the $DC$ term from the final classification layer. This single-query design substantially reduces the number of token interactions, making the decoder computationally efficient and suitable for real-time execution on vehicular hardware.
III-E3 End-to-End Computational Cost
For a single video clip and spatiotemporal view, total forward-pass cost is given by:
$\small F_{\text{total}}=F_{\text{enc}}+F_{\text{probe}}.$ (6)
If $V$ different spatiotemporal views (e.g., crops or augmentations) are processed per clip, the total cost scales linearly with $V$. Given a device with computational throughput $\phi$ (in FLOPs/s) and per-view input/output latency $t_{\text{I/O}}$, the total inference time can be approximated as:
$\small t_{\text{infer}}={F_{\text{total}}}/{\phi}+V\,t_{\text{I/O}}.$ (7)
Hence, the encoder dominates the total computational cost, while the attentive probe and classifier introduce only minor overhead ($\mathcal{O}(D^{2}+DC)$). This computational asymmetry justifies deploying the encoder at the RSU (with higher compute resources) and the lightweight decoder on the vehicle side.
IV Performance Evaluations
IV-A Evaluation Setup and Parameters
In the simulation setup, as mentioned in Section III-A1, four different cameras are used to collect consecutive frames at 50 ms intervals, which are then used to generate video clips at 20 FPS. In total, 500 traffic scenario clips have been collected, with 115 including collisions and 385 consisting of safe driving [^1] Urban Mobility Research Video Dataset (Generated with Quanser Interactive Lab), Available at: https://tinyurl.com/Uottawa-vjepa. Each frame of the videos has a 2048x2048 dimension, and the durations vary depending on the simulation runtime. The simulation automatically finalizes itself once a collision occurs. Each video’s total lengths are adjusted to a maximum of 64 frames by removing from the beginning to make it adaptable for V-JEPA training parameters. For various post-processing methods, explained in Section III-A2, a previously trained YOLOv11 model is employed from [11].
For the V-JEPA model, a ViT-Huge (ViT-H) based encoder model, which is trained using masked modeling self-supervised learning, as explained in Section III-B1, is employed directly from its repository [3], having been pretrained with VideoMix2M beforehand. ViT-H requires 384x384 video frames, so each clip, before forwarding to the V-JEPA encoder, is resized to the proper shape. Moreover, the patch size in its training is a 16x16 pixel block, and the embedding output dimension of ViT-H is a 1280-sized vector, which is used in SemComm. For fine-tuning to a specific task, as explained in Section III-B2, the K400 attentive probe is utilized. After video clips are forwarded to the frozen encoder of V-JEPA, the K400 attentive probe is trained with 40 epochs and an 8-batch size. The output layer of the attentive probe consists of two nodes: Safe Driving and Collision Probability. During the training, the AdamW optimizer is utilized with 0.001 initial learning rate of 0.001. All training parameters can be seen in Table I.
TABLE I: Performance Evaluation Setup
Total Number of Video Clips 500
#Safe/Collision Scenario 385/115
#Maximum Frame (N), FPS 64, 20
V-Jepa Encoder ViT-H
Attentive Probe K400
Input Frame Size (WxH) 384x384
Patch Size (pXp) 16x16
Embedding Dimension (D) 1280
Output of Attentive Probe 2
#Epochs, Optimizer 40, AdamW
Initial Learning Rate 0.001
IV-B Numerical Results
In this part, performance evaluation and numerical results for both the communication and the V-JEPA model are analyzed. In the communication part, there are two criteria: average compression ratio and transmission latency, as shown in Figure4. Note that we set SnR for BPSK and QAM16 to 12 dB and 22 dB, which are proven to be most efficient values in our previous work [12]
Figure 4: Results for Average Compression Ratio and Average V2X Latency For Each Embedding Format
The above part of Figure4 shows the compression ratio, while the bottom plot illustrates the average V2X transmission ratio. Raw videos collected in various experiments have variable durations, impacting their total size. In terms of compression ratio, we categorized videos into lower, standard, and higher duration categories. We calculated the compression ratios for each category’s embedding encodings (FP32, FP16, INT8), as explained in III-D.
In transmission latency, for BPSK and QAM16 modulation schemes, we set uplink data rate to 10 MHz and 20 MHz, and calculate the average transmission latency for FP32, FP16, and INT8 embedding encodings. Once the uplink rate increases, average transmission latency is reduced for both modulation schemes, and the results prove that QAM16 has much lower latency for BPKS, which makes it more suitable for real-time.
On the other hand, the evaluation of the V-JEPA model on the traffic collision avoidance is performed in two steps. In the first step, we determine which post-processing methods (Section III-A2) yield the optimal performance for V-JEPA model collision estimation. Then, we select the best postprocessing and adjust the interval before the collision occurs to see how early and accurately it can be detected. Results for different post-processing methods and original clips can be seen in Figure5 via confusion matrices.
Figure 5: Results for Different Post-Processing Techniques
Figure5 presents 4 confusion matrices for V-JEPA models; No post-processing (upper left), heatmap (upper right), binary masked (bottom left), and heatmap+binary masked (bottom right). For collision scenarios, each video clip was trimmed to exclude 8 frames depicting or directly preceding the collision event. Model achieves 76% F1 scores without post-processing, 80% heatmap, 84% binary masked, and 81.81% heatmap+binary masked. Heatmap+binary masked (3) has a lower false positive than single binary masked (6), but binary masked has the highest F1 scores due to its low false negatives. This indicates that YOLO-based vehicle highlighting compromises performance in false negative cases, despite most collisions being detected properly. In case 2, we examine how the number of removed frames before collisions affects performance (see Table II).
TABLE II: Results for Different Previous-Frames Gap Before Collisions
12-Frames
Before Collision
8-Frames
Before Collision
4-Frames
Before Collision
Accuracy 90% 92% 90%
Precision 60.9% 91.3% 82.6%
Recall 93.3% 77.8% 76.0%
F1-score 73.8% 84.0% 79.1%
V Conclusions
This paper presented a V2X SemComm framework for proactive collision prediction using the V-JEPA. RSUs serve as semantic encoders, generating predictive spatiotemporal embeddings from video data captured via the QLabs digital twin. Post-processed clips highlighting critical regions were used with pretrained V-JEPA encoders and K400 attentive probes to convert frame sequences into compact vectors capturing essential motion cues, achieving an 84% F1 score. By transmitting low-dimensional embeddings instead of raw video, the framework reduces communication payloads by up to four orders of magnitude, enabling bandwidth-efficient and low-latency semantic message exchange. Future work will explore adaptive semantic compression and cross-vehicle embedding fusion to enhance cooperative perception and scalability in next-generation V2X systems.
Acknowledgment
This work is supported in part by the MITACS Accelerate Program, in part by the NSERC CREATE TRAVERSAL program, and in part by the Ontario Research Fund-Research Excellence program under grant number ORF-RE012-026. The authors would like to thank Quanser for their support in the generation of the traffic data via Quanser Interactive Lab (https://www.quanser.com/products/quanser-interactive-labs/).

### v2

@@
 In this work, we propose a semantic V2X framework in which RSU-mounted cameras generate spatiotemporal semantic embeddings of future frames using the Video Joint Embedding Predictive Architecture (V-JEPA).
+Afterwards, These embedding future frames are utilized to estimate the likelihood collision for alert the prevention system at vehicle side.
 To evaluate the system, we construct a digital twin of an urban traffic environment enabling the generation of diverse traffic scenarios with both safe and collision events.
-These embeddings of the future frame, extracted from V-JEPA, capture task-relevant traffic dynamics and are transmitted via V2X links to vehicles, where a lightweight attentive probe and classifier decode them to predict imminent collisions.
-By transmitting only semantic embeddings instead of raw frames, the proposed system significantly reduces communication overhead while maintaining predictive accuracy.
-Experimental results demonstrate that the framework with an appropriate processing method achieves a 10% F1-score improvement for collision prediction while reducing transmission requirements by four orders of magnitude compared to raw video.
+The embeddings of the future frame, extracted from V-JEPA, capture task-relevant traffic dynamics and are transmitted via V2X links to vehicles, where a lightweight classifier decode them to predict imminent collisions.
+By transmitting only semantic embeddings instead of raw frames, the proposed system significantly reduces communication payload while maintaining predictive accuracy.
+Experimental results demonstrate that the proposed framework with an appropriate processing method achieves 92% accuracy and 8% F1-score improvement for collision prediction.
+It also reduces transmission payload by five orders of magnitude compared to raw video fulfilling the latency constraints for safe driving.
 This validates the potential of semantic V2X communication to enable cooperative, real-time collision prediction in ITS.
 I Introduction
-Traffic collisions remain one of the most severe challenges in Intelligent Transportation Systems (ITS), often leading to congestion, property damage, injuries, and fatalities.
-Proactive collision prediction is essential to mitigate these risks by enabling early warnings and preventive actions [9].
-Traditional approaches that rely on transmitting raw video [16] or high-dimensional sensory data from roadside units (RSUs) to vehicles are not feasible under V2X constraints, where both bandwidth and latency are limited.
-This calls for communication-efficient, task-oriented methods that transmit only the most relevant information to support predictive safety applications.
+Achieving proactive collision prediction is essential for enhancing road safety through early warnings and coordinated preventive actions in Intelligent Transportation Systems (ITS) [1].
+However, conventional approaches that transmit raw video streams [2] or high-dimensional sensor data from roadside units (RSUs) to vehicles are impractical under vehicular communication (V2X) constraints due to limited bandwidth and stringent latency requirements.
+These limitations highlight the need for communication-efficient, task-oriented frameworks that convey only semantically relevant information to support predictive safety functions in connected and autonomous vehicles.
 Semantic communication (SemComm) addresses this by transmitting task-relevant features instead of raw sensory data [3].
@@
 The digital twin thus enables systematic video-data generation for training and evaluating predictive SemComm frameworks.
-Building on this foundation, we propose a semantic V2X collision prediction framework.
+Building on this foundation, we propose a novel semantic V2X collision prediction framework which departs from conventional perception-centric V2X approaches by performing future-oriented semantic prediction.
 RSUs equipped with cameras employ the Video Joint Embedding Predictive Architecture (V-JEPA) as the semantic encoder [4].
 V-JEPA does not merely summarize observed frames; it predicts the embeddings of future frames, capturing the evolving motion patterns and vehicle interactions that lead to collisions.
-These predicted embeddings are transmitted to vehicles via V2X links, where an attentive probe and classifier act as semantic decoders, refining the embeddings to predict whether a collision is imminent.
+These predicted embeddings are transmitted to vehicles via V2X links, where a classifier acts as a semantic decoder, refining the embeddings to predict whether a collision is imminent.
 By focusing on predictive embeddings, the system moves beyond detection toward true anticipation of dangerous events.
 The main contributions of this work are summarized as follows:
-  • We construct a high-quality video dataset using a digital twin of an urban traffic environment (QLabs).
-The dataset includes diverse scenarios and is labeled into two categories, collision events and safe driving, enabling systematic evaluation of predictive collision avoidance.
-  • We investigate multiple post-processing techniques (heatmaps, binary masking, and hybrid methods) to enhance task-relevant features in traffic videos, and analyze their impact on the representation quality and predictive performance of the V-JEPA framework.
-  • We employ V-JEPA as the semantic encoder at RSUs to extract low-dimensional spatiotemporal embeddings of future frame from video clips.
-These embeddings serve as semantic messages transmitted via V2X links, and are decoded by an attentive probe with cross-attention at the vehicle side to focus on high-risk regions and motion patterns for accurate collision prediction.
-This work demonstrates that combining digital twins, semantic communication, and spatiotemporal embeddings can enable predictive collision avoidance in vehicular networks, achieving both high accuracy and significant communication efficiency.
+  • A high-quality video dataset is generated using a QLabs digital twin, featuring diverse urban scenarios labeled as collision or safe, supporting systematic evaluation of predictive collision avoidance.
+  • We evaluate post-processing methods such as heatmaps, binary masks, and hybrids to enhance task-relevant features and assess their effect on V-JEPA’s representation quality and predictive accuracy.
+  • V-JEPA is employed to encode future-frame embeddings at RSUs, which are transmitted via V2X and decoded by a lightweight classifier in vehicles to identify high-risk regions for collision prediction.
+This work demonstrates that combining digital twins, SemComm, and spatiotemporal embeddings can enable predictive collision avoidance in vehicular networks, achieving 92% accuracy and 8% F1-score improvement in collision prediction with five orders of magnitude reduction in transmission (Tx) payload, supporting real-time latency constraint V2X communication.
 II Related Works
-Collision prediction and avoidance have been studied from multiple perspectives.
-The authors in [8] propose a deep learning-based approach using convolutional neural networks (CNNs) to analyze surveillance video for traffic congestion detection and accident monitoring, aiming to overcome the limitations of manual observation and traditional machine learning models.
-Similarly, [2] employs ensemble learning methods such as Random Forest, XGBoost, L-GBM, and CatBoost combined with Shapley Additive Explanations (SHAP) to predict accident severity and identify contributing factors, providing interpretable insights into road safety analysis.
-These works, however, largely focus on post-hoc analysis of accidents rather than proactive prediction of imminent collisions.
+Collision prediction and avoidance have been explored from various perspectives.
+In [5], a deep learning model based on convolutional neural networks (CNNs) analyzes surveillance video for traffic congestion and accident detection, addressing the limits of manual observation and conventional learning methods.
+Similarly, the work in [6] uses ensemble models (Random Forest, XGBoost, L-GBM, CatBoost) with SHAP-based interpretability to assess accident severity and contributing factors.
+However, these studies mainly emphasize post-event analysis rather than proactive prediction of collisions.
 Big data frameworks have also been leveraged in [7], where large-scale UK traffic datasets were processed using Apache Spark with Decision Trees, Random Forest, and Naïve Bayes to predict accident severity under diverse environmental conditions.
@@
 While these approaches improve interpretability, they rely heavily on temporal correlations in data streams rather than predictive spatiotemporal embeddings tailored for proactive vehicular safety.
-YOLO-based detection frameworks are also gaining traction.
-In [10], YOLOv8 and YOLO11x models were trained on 15,000 traffic images for real-time accident and speeding-vehicle detection, introducing persistent multi-frame verification and human-in-the-loop validation.
-For proactive collision avoidance, [5] integrates CARLA simulations with YOLOv7-based object detection to anticipate collisions in connected and autonomous vehicle scenarios.
-Similarly, [17] enhances small-object detection and near-miss prediction by combining YOLOv7 with instance segmentation, Inverse Perspective Mapping (IPM), StrongSORT tracking, and Kalman filtering.
-These YOLO-based methods are effective in object-level detection and trajectory forecasting but still require transmitting large amounts of visual data for centralized processing.
-In contrast, this work proposes a V2X SemComm framework where RSUs generate predictive spatiotemporal embeddings using V-JEPA.
-Unlike CNN- or YOLO-based methods that detect current events, V-JEPA forecasts embeddings of future frames, anticipating motion patterns that may lead to collisions.
-Instead of transmitting raw video, embeddings are sent as semantic messages to vehicles, which decode them with a lightweight attentive probe and classifier.
-This design enables real-time, bandwidth-efficient collision prediction, and, to our knowledge, is the first to integrate digital twin–based data generation, predictive embeddings, and semantic V2X communication.
+YOLO-based detection frameworks have gained attention for traffic analysis.
+In [11], YOLOv8 and YOLO11x models are trained on 15,000 traffic images for real-time accident and speeding-vehicle detection.
+For proactive collision avoidance, the work in [12] integrates CARLA simulations with YOLOv7-based object detection, while the study in [13] enhances small-object detection and near-miss prediction through instance segmentation, IPM, and tracking modules.
+Although effective for object-level detection, these methods still rely on transmitting large volumes of visual data for centralized processing.
+In contrast, this work introduces a V2X semantic communication framework where RSUs use V-JEPA to generate predictive spatiotemporal embeddings of future frames, enabling anticipation of collision-prone motion patterns.
+Rather than transmitting raw video, semantic embeddings are shared with vehicles for decoding via a lightweight classifier, achieving real-time and bandwidth-efficient collision prediction.
 III Methodology
@@
 Instead of reconstructing raw frames, V-JEPA generates compact spatiotemporal embeddings of future frames, thereby anticipating anomalous motion patterns that precede collisions.
-These predicted embeddings are transmitted as semantic messages over V2X links to vehicles, where an attentive probe and classifier decode them to determine whether a collision is imminent or the driving condition is safe.
-By transmitting embeddings rather than raw video, the system reduces both communication overhead and computational workload while preserving predictive accuracy.
+These predicted embeddings are transmitted as semantic messages over V2X links to vehicles, where a classifier decode them to determine whether a collision is imminent or the driving condition is safe.
+By transmitting embeddings rather than raw video, the system reduces both communication payload and computational workload while preserving predictive accuracy.
 Figure 1: Data Collection Using Digital-Twin (Quanser Interactive Lab)
@@
 The Qlabs environment is programmed to generate both safe driving and collision events, observed from four RSU-mounted surveillance cameras covering different regions of the environment.
-Each frame is sampled every 50 ms, producing 20 FPS (Frames per second) video clips of varying lengths depending on scenario duration.
+Simulation provides 4 different type of regions: 4-way junction, 3-way junction, side roads and roundabout.
+There are total 500 video clips and each frame is sampled every 50 ms, producing 20 FPS (Frames per second) of varying lengths depending on scenario duration.
 This allows for reproducible traffic scene generation and controlled experimentation for predictive safety analysis.
@@
 The V-JEPA model [4] is deployed at RSUs to generate predictive semantic embeddings.
-Unlike conventional predictive models, V-JEPA forecasts embeddings of masked future frame regions rather than reconstructing pixels, thereby focusing on task-relevant semantics for downstream prediction.
+Unlike conventional predictive models, V-JEPA forecasts embeddings of masked future frame rather than reconstructing pixels, thereby focusing on task-relevant semantics for downstream prediction.
 III-B1 V-JEPA Pretraining with Masked Modeling
 As shown in Figure 3, V-JEPA is pretrained in a self-supervised fashion using masked prediction.
-A video clip is divided into non-overlapping spatiotemporal patches.
+A video clip is divided into non-overlapping spatiotemporal patches, where each frame of spatial size $H\times W$ is partitioned into $P=\tfrac{H}{p}\times\tfrac{W}{p}$ patches based on a patch size of $p\times p$.
 The target encoder processes the full unmasked clip, while the context encoder encodes a masked version.
-A predictor network generates embeddings for the masked tokens, which are trained against target encoder embeddings using an $L_{1}$ loss.
-This allows the model to learn predictive spatiotemporal representations in the embedding space.
-Figure 3: V-JEPA Pretraining & Fine-Tuning for Collision Prediction and V2X Semantic Communication
+Each patch is projected into a $D$-dimensional embedding vector, resulting in a per-frame embedding tensor of size $P\times D$.
+A predictor network then generates embeddings for the masked tokens, which are trained against the corresponding target encoder embeddings using an $L_{1}$ loss.
+This enables the model to learn predictive spatiotemporal representations directly in the embedding space rather than at the pixel level.
+Figure 3: V-JEPA Pretraining and Fine-Tuning for Collision Prediction
 III-B2 Fine-Tuning with Attentive Probe
-For downstream collision prediction, the pretrained V-JEPA encoder takes as input a video clip of dimension $N\times W\times H$, where $N$ is the number of frames, and $W$ and $H$ denote frame width and height.
+For downstream collision prediction, the pretrained V-JEPA encoder takes as input a video clip of dimension $N\times W\times H$, where $N$ is the number of frames in the clip.
 The clip is flattened into a sequence of non-overlapping spatiotemporal patches, producing token embeddings of shape $N\times P\times D$, where $P$ is the number of patches per frame and $D$ is the embedding dimension.
 These frozen encoder parameters preserve the rich self-supervised features learned during pretraining.
-On top of the encoder outputs, an attentive probe applies a query-driven cross-attention mechanism to pool the tokens into a single compact vector $1\times D$, selectively aggregating information from the most relevant spatial and temporal regions.
+On top of the encoder outputs, The attentive probe employs a query-driven cross-attention mechanism, where the query vector acts as a learnable global token that attends to the most informative spatial and temporal regions in the video.
+Formally, let the encoder output token embeddings be $\mathbf{Z}\in\mathbb{R}^{(N\cdot P)\times D}$.
+The attentive probe introduces a learnable query matrix $\mathbf{Q}\in\mathbb{R}^{Q\times D}$, with $Q=1$ in this setup, to perform cross-attention as follows:
+$\small\mathbf{z}_{\text{attn}}=\text{Attn}(\mathbf{Q},\mathbf{Z})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{Z}^{\top}}{\sqrt{D}}\right)\mathbf{Z},$ (1)
+where $\mathbf{Q}$ serves as the query, and $\mathbf{Z}$ provides both keys and values.
+The resulting weighted aggregation yields a single compact embedding $\mathbf{z}_{\text{attn}}\in\mathbb{R}^{1\times D}$, which captures the most salient spatiotemporal dynamics relevant to collision prediction.
 This pooled embedding is then passed to a lightweight classifier to infer one of two states: collision or safe driving.
-By transforming raw video data ($N\times W\times H$) into a compact embedding $1\times D$, the framework enables efficient semantic transmission while maintaining predictive accuracy for V2X communication.
+By transforming raw video data ($N\times W\times H$) into a compact embedding $1\times D$, the framework enables efficient semantic Tx while maintaining predictive accuracy for V2X communication.
 III-C Semantic Decoder at Vehicle
-At the vehicle side, the received embeddings are directly processed by the semantic decoder, which consists of an attentive probe and a lightweight classifier.
-Vehicles receive embeddings from RSUs, refine them through query-driven cross-attention, concentrating on the most critical spatial and temporal cues that may indicate anomalous motion or collision risk.
+At the vehicle side, the received embeddings are directly processed by the semantic decoder, which consists of a lightweight classifier.
+Vehicles receive embeddings from RSUs, that are refined through query-driven cross-attention, concentrating on the most critical spatial and temporal cues that may indicate anomalous motion or collision risk.
 The resulting compact representation is then fed into a linear classifier to produce a binary decision: collision or safe driving.
@@
 III-D Transmission Efficiency in V2X Semantic Communication
-A major advantage of the proposed framework lies in its ability to drastically reduce transmission payloads compared to conventional raw video communication.
+A major advantage of the proposed framework lies in its ability to drastically reduce Tx payloads compared to conventional raw video communication.
 Let each video clip contain $N$ frames of original spatial size $H_{o}\times W_{o}$ with 3 RGB channels, where $(H_{o},W_{o})$ denote the preprocessed frame dimensions.
@@
 Each clip produces embeddings of size $N\times P\times D$, where $P$ is the number of spatiotemporal patches per frame and $D$ is the embedding dimension.
-The corresponding payload is:
-$\small S_{\text{sem}}=N\cdot P\cdot D\cdot b\quad\text{bytes},$ (2)
-where $b$ denotes the number of bytes per element ($b=4$ for FP32, $b=2$ for FP16, $b=1$ for INT8).
-The semantic compression ratio is then: $\mathcal{R}={S_{\text{raw}}}/{S_{\text{sem}}}$.
-Assuming $N=64$, $H_{o}=W_{o}=2048$, and a post-processed input of $384\times 384$ with patch size $16\times 16$ ($P=\tfrac{384}{16}\times\tfrac{384}{16}=576$) and embedding dimension $D=1280$, the corresponding payloads are:
-  • Raw video (RGB): 0.81 GB per clip.
-  • Embeddings (FP32): 0.74 MB ($\mathcal{R}\approx 1095\times$).
-  • Embeddings (FP16): 0.37 MB ($\mathcal{R}\approx 2190\times$).
-  • Embeddings (INT8): 0.18 MB ($\mathcal{R}\approx 4380\times$).
-These results demonstrate that transmitting semantic embeddings instead of raw video frames achieves up to four orders of magnitude reduction in communication payload.
-This enables low-latency, bandwidth-efficient semantic message exchange between RSUs and vehicles, ensuring timely and reliable collision prediction in V2X networks.
+Then, an attentive probe aggregates these tokens into a single compact representation of size $1\times D$ using a query-driven cross-attention mechanism.
+This process selectively emphasizes the most informative spatial and temporal regions, resulting in a lightweight semantic vector suitable for V2X transmission.
+The payload corresponding to the transmitted semantic embedding is therefore:
+$\small S_{\text{sem}}=1\times D\times b\quad\text{bytes},$ (3)
+where $b$ denotes the number of bytes per element ($b=4$ for FP32, $b=2$ for FP16, $b=1$ for INT8). The semantic compression ratio can be defined as $\mathcal{R}={S_{\text{raw}}}/{S_{\text{sem}}}.$
 III-E Complexity Analysis of Frozen Encoder with Attentive Probe
@@
 where $C$ denotes the number of output classes (e.g., $C=2$ for collision and safe driving).
-However, in practice, the encoder output is frozen and the probe operates with a single query vector rather than performing full self-attention over all $L$ tokens.
+However, in practice, the encoder output is frozen and the probe operates with a single query vector rather than performing full self-attention over all $L$ tokens to capture a unified global representation.
 Consequently, the cost associated with key and value projections ($\mathcal{O}(LD^{2})$) is computed once during encoding and not repeated during probing.
 The probe therefore performs only lightweight attention ($\mathcal{O}(LD)$) and projection operations ($\mathcal{O}(D^{2})$), resulting in an overall complexity that scales as:
-$\small\mathcal{O}(D^{2}+DC),\small$ (5)
+$\mathcal{O}(D^{2}+DC),\small$ (6)
 where the $D^{2}$ term arises from the linear transformations within the probe and the $DC$ term from the final classification layer.
@@
 If $V$ different spatiotemporal views (e.g., crops or augmentations) are processed per clip, the total cost scales linearly with $V$.
-Given a device with computational throughput $\phi$ (in FLOPs/s) and per-view input/output latency $t_{\text{I/O}}$, the total inference time can be approximated as:
-$\small t_{\text{infer}}={F_{\text{total}}}/{\phi}+V\,t_{\text{I/O}}.$ (7)
-Hence, the encoder dominates the total computational cost, while the attentive probe and classifier introduce only minor overhead ($\mathcal{O}(D^{2}+DC)$).
+Given a device with computational throughput $\phi$ (in FLOPs/s) and per-view input/output latency $t_{\text{I/O}}$, the total inference time can be approximated as $t_{\text{infer}}={F_{\text{total}}}/{\phi}+V\,t_{\text{I/O}}.$ Hence, the encoder dominates the total computational cost, while the attentive probe and classifier introduce only minor overhead ($\mathcal{O}(D^{2}+DC)$).
 This computational asymmetry justifies deploying the encoder at the RSU (with higher compute resources) and the lightweight decoder on the vehicle side.
-IV Performance Evaluations
+IV Performance Evaluation
 IV-A Evaluation Setup and Parameters
-In the simulation setup, as mentioned in Section III-A1, four different cameras are used to collect consecutive frames at 50 ms intervals, which are then used to generate video clips at 20 FPS.
-In total, 500 traffic scenario clips have been collected, with 115 including collisions and 385 consisting of safe driving [^1] Urban Mobility Research Video Dataset (Generated with Quanser Interactive Lab), Available at: https://tinyurl.com/Uottawa-vjepa.
-Each frame of the videos has a 2048x2048 dimension, and the durations vary depending on the simulation runtime.
+As mentioned in Section III-A1, four different cameras are used in the QLabs platform to collect consecutive frames at 50 ms intervals, which are then used to generate video clips at 20 FPS.
+In total, 500 traffic scenario clips have been collected, with 115 including collisions and 385 consisting of safe driving [^1] Urban Mobility Research Video Dataset, Available at: https://tinyurl.com/Uottawa-vjepa.
+Generated with Quanser Interactive Lab: (https://www.quanser.com/products/quanser-interactive-labs/).
+The clips’ durations vary depending on the simulation runtime.
 The simulation automatically finalizes itself once a collision occurs.
-Each video’s total lengths are adjusted to a maximum of 64 frames by removing from the beginning to make it adaptable for V-JEPA training parameters.
+The long clips’ lengths are adjusted to a maximum of 64 frames by removing from the beginning to make it adaptable for V-JEPA training parameters.
 For various post-processing methods, explained in Section III-A2, a previously trained YOLOv11 model is employed from [16].
-For the V-JEPA model, a ViT-Huge (ViT-H) based encoder model, which is trained using masked modeling self-supervised learning, as explained in Section III-B1, is employed directly from its repository [3], having been pretrained with VideoMix2M beforehand.
-ViT-H requires 384x384 video frames, so each clip, before forwarding to the V-JEPA encoder, is resized to the proper shape.
-Moreover, the patch size in its training is a 16x16 pixel block, and the embedding output dimension of ViT-H is a 1280-sized vector, which is used in SemComm.
-For fine-tuning to a specific task, as explained in Section III-B2, the K400 attentive probe is utilized.
-After video clips are forwarded to the frozen encoder of V-JEPA, the K400 attentive probe is trained with 40 epochs and an 8-batch size.
-The output layer of the attentive probe consists of two nodes: Safe Driving and Collision Probability.
-During the training, the AdamW optimizer is utilized with 0.001 initial learning rate of 0.001.
-All training parameters can be seen in Table I.
-TABLE I: Performance Evaluation Setup
+For the V-JEPA model, a ViT-Huge (ViT-H) encoder pretrained with VideoMix2M via masked modeling self-supervised learning [4] is employed, as described in Section III-B1.
+Each video clip is resized to 384×384 before encoding, with a patch size of 16×16 and an embedding dimension of 1280 used in the SemComm framework.
+For task-specific fine-tuning (Section III-B2), the K400 attentive probe is trained for 40 epochs with a batch size of 8, using the AdamW optimizer and an initial learning rate of 0.001.
+The output layer includes two nodes, namely the Safe Driving and Collision Probability, and the training parameters are summarized in Table I.
+TABLE I: Training Setup and Parameters
 Total Number of Video Clips 500
 #Safe/Collision Scenario 385/115
-#Maximum Frame (N), FPS 64, 20
+Max. Video Length (N), FPS 64 Frames, 20
 V-Jepa Encoder ViT-H
 Attentive Probe K400
+Original Frame Size $(W_{o}\times H_{o})$ 2048x2048
 Input Frame Size (WxH) 384x384
@@
 Embedding Dimension (D) 1280
-Output of Attentive Probe 2
+Output of Classifier 2
 #Epochs, Optimizer 40, AdamW
 Initial Learning Rate 0.001
+Modulation Schemes/SNR BPSK/12dB, QAM-16/22dB [17]
+Encoding Format INT8, FP16, FP32
+Bandwidth 20 MHz
 IV-B Numerical Results
-In this part, performance evaluation and numerical results for both the communication and the V-JEPA model are analyzed.
-In the communication part, there are two criteria: average compression ratio and transmission latency, as shown in Figure4.
-Note that we set SnR for BPSK and QAM16 to 12 dB and 22 dB, which are proven to be most efficient values in our previous work [12]
-Figure 4: Results for Average Compression Ratio and Average V2X Latency For Each Embedding Format
-The above part of Figure4 shows the compression ratio, while the bottom plot illustrates the average V2X transmission ratio.
-Raw videos collected in various experiments have variable durations, impacting their total size.
-In terms of compression ratio, we categorized videos into lower, standard, and higher duration categories.
-We calculated the compression ratios for each category’s embedding encodings (FP32, FP16, INT8), as explained in III-D.
-In transmission latency, for BPSK and QAM16 modulation schemes, we set uplink data rate to 10 MHz and 20 MHz, and calculate the average transmission latency for FP32, FP16, and INT8 embedding encodings.
-Once the uplink rate increases, average transmission latency is reduced for both modulation schemes, and the results prove that QAM16 has much lower latency for BPKS, which makes it more suitable for real-time.
-On the other hand, the evaluation of the V-JEPA model on the traffic collision avoidance is performed in two steps.
-In the first step, we determine which post-processing methods (Section III-A2) yield the optimal performance for V-JEPA model collision estimation.
-Then, we select the best postprocessing and adjust the interval before the collision occurs to see how early and accurately it can be detected.
-Results for different post-processing methods and original clips can be seen in Figure5 via confusion matrices.
+From a communication perspective, two key metrics are analyzed: compression ratio and Tx latency (Figure 4).
+Figure 4: Average compression ratio for different video lengths and Tx latency for different modulation schemes
+The left portion of Figure 4 illustrates the achieved compression ratios, while the right portion depicts the corresponding Tx latencies.
+The raw videos obtained from different experiments vary in duration, resulting in different overall data sizes.
+To assess compression efficiency, the videos are categorized into two duration groups: normal-length and large-length clips.
+For each group, compression ratios are computed across the three encoding formats (FP32, FP16, and INT8), as described in Section III-D.
+In both cases, INT8 achieves the highest compression ratio, exceeding $10^{5}$, followed by FP16 and FP32.
+Large-length clips (64 frames) yield higher ratios than normal-length clips (33 frames), reflecting reduced transmission payload with longer videos.
+For example, assuming $N=64$, $H_{o}=W_{o}=2048$, the raw video size is approximately $64\times 2048\times 2048\times 3=0.81$ GB, while the transmitted semantic embeddings require only $1\times 1280\times b=0.005$ MB, $0.0025$ MB, and $0.0012$ MB for FP32, FP16, and INT8, respectively, corresponding to compression gains of $1.6\times 10^{5}$, $3.2\times 10^{5}$, and $6.4\times 10^{5}$.
+To evaluate latency and Tx data rate, the SNR values for BPSK and QAM16 are set to 12 dB and 22 dB, respectively, as identified optimal previously [17].
+QAM16 achieves the lowest Tx latency across all encodings, 0.27, 0.13, and 0.06 ms for FP32, FP16, and INT8,while BPSK yields 0.50, 0.25, and 0.12 ms under the same conditions.
+All values remain well below the 5 ms V2X threshold, confirming the framework’s suitability for real-time, low-latency vehicular communication.
+The evaluation of the V-JEPA model for traffic collision avoidance is conducted in two stages.
+First, various post-processing methods (Section III-A2) are analyzed to identify the configuration that provides the highest collision prediction accuracy.
+Next, the selected post-processing approach is used to adjust the frame gap preceding the collision to assess how early and reliably potential collisions can be detected.
+The resulting confusion matrices for different post-processing methods and the original clips are presented in Figure 5.
 Figure 5: Results for Different Post-Processing Techniques
-Figure5 presents 4 confusion matrices for V-JEPA models; No post-processing (upper left), heatmap (upper right), binary masked (bottom left), and heatmap+binary masked (bottom right).
+Figure5 presents the confusion matrices for no post-processing (upper left), heatmap (upper right), binary masked (bottom left), and heatmap+binary masked (bottom right).
 For collision scenarios, each video clip was trimmed to exclude 8 frames depicting or directly preceding the collision event.
 Model achieves 76% F1 scores without post-processing, 80% heatmap, 84% binary masked, and 81.81% heatmap+binary masked.
-Heatmap+binary masked (3) has a lower false positive than single binary masked (6), but binary masked has the highest F1 scores due to its low false negatives.
+Heatmap+binary masked has a lower false positive than binary masked, but binary masked has the highest F1 scores due to its low false negatives.
 This indicates that YOLO-based vehicle highlighting compromises performance in false negative cases, despite most collisions being detected properly.
-In case 2, we examine how the number of removed frames before collisions affects performance (see Table II).
-TABLE II: Results for Different Previous-Frames Gap Before Collisions
+In the second step, we examine how the number of removed frames (frame-gap) before collisions affects collision prediction performance of V-JEPA Model (see Table II).
+TABLE II: Results for Different Previous-Frames Gap
 12-Frames
@@
 F1-score 73.8% 84.0% 79.1%
+The number of removed frames before the collisions indicates how early we can detect the collision.
+If the frame-gap before collision increases, the model tries to estimate the likelihood of future collision earlier.
+To achieve better results, we set the frame-gap to 12 (left), 8 (middle), and 4 (right) with binary mapped clips, resulting in 73.68%, 84%, and 79% F1 scores, respectively.
+With 12 frames removed before the collision, the model yields highly false negative values because many collisions occur during cornering, and a longer frame cut results in clips ending before cornering begins.
+The 4-frame gap has a slightly lower F1 score than 8-frame gap before colliding because the model is already cornering and 4-frame removing used longer clips, making the model’s work challenging.
 V Conclusions
-This paper presented a V2X SemComm framework for proactive collision prediction using the V-JEPA.
+This paper has presented a V2X SemComm framework for proactive collision prediction using the V-JEPA.
 RSUs serve as semantic encoders, generating predictive spatiotemporal embeddings from video data captured via the QLabs digital twin.
-Post-processed clips highlighting critical regions were used with pretrained V-JEPA encoders and K400 attentive probes to convert frame sequences into compact vectors capturing essential motion cues, achieving an 84% F1 score.
-By transmitting low-dimensional embeddings instead of raw video, the framework reduces communication payloads by up to four orders of magnitude, enabling bandwidth-efficient and low-latency semantic message exchange.
-Future work will explore adaptive semantic compression and cross-vehicle embedding fusion to enhance cooperative perception and scalability in next-generation V2X systems.
+Post-processed clips highlighting critical regions were used with pretrained V-JEPA encoders and K400 attentive probes to convert frame sequences into compact vectors capturing essential motion cues, achieving 92% accuracy and 8% F1-score improvement in collision prediction.
+By transmitting low-dimensional embeddings instead of raw video, the framework reduces communication payloads by up to five orders of magnitude, enabling bandwidth-efficient and low-latency semantic message exchange.
+We are currently exploring cross-vehicle embedding fusion to enhance cooperative perception and scalability in next-generation V2X systems while working on the real dataset in more complex scenarios.
 Acknowledgment
 This work is supported in part by the MITACS Accelerate Program, in part by the NSERC CREATE TRAVERSAL program, and in part by the Ontario Research Fund-Research Excellence program under grant number ORF-RE012-026.
-The authors would like to thank Quanser for their support in the generation of the traffic data via Quanser Interactive Lab (https://www.quanser.com/products/quanser-interactive-labs/).
+The authors would like to thank Quanser for their support in the generation of the traffic data via Quanser Interactive Lab.